Question 162554: 9. (4, -10), (1, -1)
x y 1 x y
4 -10 1 4 -10
1 -1 1 1 -1
(x)(-10)(1)+(y)(1)(1)+(1)(4)(-1)+(1)(-10)(1)-(-1)(1)(x)-(1)(4)(y)=
-10x+y+(-4)-(-10)-(-x)-4y=
-9x-3y+6=0
-9x-3y=-6
-3y+(-9x)=-6
-3y=-6+9x
-y=2+3x
Y=-3x-2
Can you find my sign error? Im sure im just overlooking something,but I can't find it. Thank you so much!

checkley77 was nice enough to try to help me,but im still confused. I see that instead of (x)(-10)(1)+(y)(1)(1)+(1)(4)(-1)+(1)(-10)(1)-(-1)(1)(x)-(1)(4)(y)=
I needed to have (x)(-10)(1)+(y)(1)(1)+(1)(4)(-1)-(1)(-10)(1)-(-1)(1)(x)-(1)(4)(y)=
But I want to know if that effects my answer? Is Y=-3x-2 correct?
: Question 162554: 9. (4, -10), (1, -1)
x y 1 x y
4 -10 1 4 -10
1 -1 1 1 -1
(x)(-10)(1)+(y)(1)(1)+(1)(4)(-1)+(1)(-10)(1)-(-1)(1)(x)-(1)(4)(y)=
-10x+y+(-4)-(-10)-(-x)-4y=
-9x-3y+6=0
-9x-3y=-6
-3y+(-9x)=-6
-3y=-6+9x
-y=2+3x
Y=-3x-2
Can you find my sign error? Im sure im just overlooking something,but I can't find it. Thank you so much!

checkley77 was nice enough to try to help me,but im still confused. I see that instead of (x)(-10)(1)+(y)(1)(1)+(1)(4)(-1)+(1)(-10)(1)-(-1)(1)(x)-(1)(4)(y)=
I needed to have (x)(-10)(1)+(y)(1)(1)+(1)(4)(-1)-(1)(-10)(1)-(-1)(1)(x)-(1)(4)(y)=
But I want to know if that effects my answer? Is Y=-3x-2 correct?

Answer by checkley77(3848)

I needed to have (x)(-10)(1)+(y)(1)(1)+(1)(4)(-1)-(1)(-10)(1)-(-1)(1)(x)-(1)(4)
-10x+y+(-4)-(-10)-(-x)-4y=0
this becomes:
-10x+y-4+10+x-4y=0
-3y-9x+6=0
-3y=9x-6
y=9x/-3-6/-3
(y)=-3x+2
But I want to know if that effects my answer? Is Y=-3x-2 correct.
I assumed you would complete the solution after the correction was provided??????